# Optimal routing and communication strategies for autonomous reconnaissance missions

### v1

@@ Abstract
We consider an autonomous reconnaissance mission where a drone has to visit several points of interest and communicate the intel back to the base. At every point of interest, the drone has the option to either send back all available info, or continue to the next point of interest and communicate at a later stage. Both choices have a chance of detection, meaning the mission fails. We wish to maximize the expected amount of information gathered by the mission. This is modeled by a routing problem in a weighted graph. We discuss the ILP formulation of this problem, show it is NP-complete, and use a genetic algorithm to find good solutions for up to ten points of interest.
@@ 1 Introduction
The role of Unmanned Aerial Vehicles (UAVs), more commonly known as drones, in society continues to become more significant every day. The civil market alone is estimated to be worth 7.2 billion USD in 2022 and this value is expected to grow to 19.2 billion USD in 2031 [2]. Applications range from agriculture over disaster response to package deliveries. But also its military use has become more relevant. Already in the Vietnam War, the US army deployed drones as a weapon [3]. When surveillance technology improved, it became clear that drones could also be used for survey missions in enemy terrain [1]. Moreover, these automated missions need not be performed by aerial systems, depending on the terrain and the characteristic of the mission, an automated ground vehicle (UGV) or an unmanned under water vehicle (UUV) may be more adept. The ongoing wars in Ukraine and Gaza have shown the importance of hybrid warfare [8]. In such warfare, the distinction between different modes of warfare (conventional warfare, cyber warfare, political warfare) tend to blurred. This makes the adversary in such war fluid and harder to predict [6]. Intelligence in physical and non-physical infrastructure is key in gaining advantage. Hence the extent to which unmanned vehicles are used for both offensive as well as reconnaissance missions is at an all-time high [10].
@@
To expand the number of operational systems while managing costs, it is desirable to deploy systems that can operate fully independently. For a reconnaissance mission, this requires a planning of the complete mission before the drone leaves for enemy territory. The setting of such a mission can be stated as follows: starting from a secure base, multiple surveillance locations need to be safely reached and the acquired information has to be brought back to the base. There are three possible ways of bringing back information. At every surveillance location, there is the possibility of transmitting information back to the base camp. Hence the first option is that the drone travels to a surveillance location, gathers the info and immediately sends it back to the base. Secondly, the drone could also store the information and go to the next surveillance location. After having obtained the information there, the totality of info could then be transmitted together at that location. Thirdly, the drone could also store information and return to the base. In this case the information is physically obtained back from the drone. However, each action in the mission carries the risk of detection – the drone could be spotted during flight, or transmissions might be intercepted. Both ways of detection give away the position of the drone, ending the mission abruptly.
@@
This paper investigates how to find the optimal strategy of these reconnaissance missions. Such an optimal strategy consists of two elements: both the route and the send strategy have to be optimal to maximize the amount of retrieved information. Hence the specific questions for which an answer is sought in this research are:
  • In which order should the different locations be traveled to?
@@
  • Where is it beneficial to make a transmission and where is it better to hold on to the gathered information?
@@ 2.2 NP-completeness
The problem lies in NP. Let $G^{\prime}=(V^{\prime},E^{\prime})$ be an undirected graph such that $0,t\in V^{\prime}$. The problem of deciding whether there is a Hamiltonian path from 00 to $t$ in $G^{\prime}$ is NP-complete, see e.g. Schrijver [9, Corollary 8.11b]. Now construct the graph $G=(V,E)$ by adding a vertex $u$ to $G^{\prime}$, connecting it only to $t$. Formally, $V=V^{\prime}\cup\{u\}$ and $E=E^{\prime}\cup\{\{t,u\}\}$. Then $G$ contains a Hamiltonian path with starting point 00 if and only if $G^{\prime}$ contains a Hamiltonian path from 00 to $t$. ∎
@@ 3.1 Moving the drone
The route of the drone will be modeled by the decision variables
@@ 4.1.1 Initialization
To initialize the algorithm, a list of random routes is generated. These routes are constructed by making a random walk of random length through the graph, starting from the base. Just as in the Linear Program, where the value of $T_{\text{max}}$ had to be guessed, the upper bound $L_{\max}$ on the length of the walk also has to be chosen here. Similarly, we also choose a lower bound $L_{\min}$. Then we sample the actual length of the random walk uniformly at random from the set $\{L_{\min},L_{\min}+1,\dots,L_{\max}-1,L_{\max}\}$. Secondly, we repeatedly pick the next vertex in the walk uniformly at random from the vertices adjacent to the current position. If the last vertex of the walk is not the base vertex, the shortest path between the end of the walk and the base is added to the walk. As a rule of thumb, when considering a graph on $n$ vertices, we picked $L_{\min}$ to be approximately equal to $n-1$ and $L_{\max}$ to $n+100$. This might seem like a pretty high upper bound, especially because realistically, reconnaissance mission do not include more than 10 surveillance locations. But even though these longer routes are not necessarily better than the shorter ones – at some point all information has been transmitted and making an additional loop doesn’t change the expected value – the final part of a route with useless and unnecessary walks can become useful after a crossover.
@@ 5.1 Adaptation of the Genetic Algorithm
Also the mutations have the same flavor as in the single drone case. First of all, there is the added random walk. This is the same mutation as described for the single drone: one of the routes is picked and a random walk is added somewhere in the walk. This continues to be a useful mutation to increase the variation in the genetic program. But it slows down the algorithm, because it requires a local search for the optimal send strategy in terms of the expected value for multiple drones. Therefore this mutation is only actualized with a low probability. If a generation consists of $\kappa$ strategies, the mutation probability is set at 2/$\kappa$.

### v2

# Optimal routing and transmission strategies for UAV reconnaissance missions with detection threats

@@ Abstract
We consider an autonomous reconnaissance mission where an Unmanned Aerial Vehicle (UAV) has to visit several points of interest and communicate the intel back to the base. At every point of interest, the UAV has the option to either send back all available info, or continue to the next point of interest and communicate at a later stage. Both choices have a chance of detection, meaning the mission fails. We wish to maximize the expected amount of information gathered by the mission. This is modelled by a routing problem in a weighted graph. We show that the problem is NP-complete, discuss an ILP formulation, and use a genetic algorithm to find good solutions for up to ten points of interest.
@@ 1 Introduction
The role of Unmanned Aerial Vehicles (UAVs), more commonly known as drones, in society continues to become more significant every day. The civil market alone is estimated to be worth 7.2 billion USD in 2022 and this value is expected to grow to 19.2 billion USD in 2031 [4]. Applications range from agriculture over disaster response to package deliveries. But also its military use has become more relevant. Already in the Vietnam War, the US army deployed drones as a weapon [7]. When surveillance technology improved, it became clear that drones could also be used for survey missions in enemy terrain [3]. Moreover, these automated missions need not be performed by aerial systems, depending on the terrain and the characteristic of the mission, an automated ground vehicle (UGV) or an unmanned under water vehicle (UUV) may be more adept. The wars in Ukraine and Gaza have once again shown the importance of hybrid warfare [16]. In such warfare, the distinction between different modes of warfare (conventional warfare, cyber warfare, political warfare) tend to blurred. This makes the adversary in such war fluid and harder to predict [12]. Intelligence in physical and non-physical infrastructure is key in gaining advantage. Hence the extent to which unmanned vehicles are used for both offensive as well as reconnaissance missions is at an all-time high [23].
@@
To expand the number of operational systems while managing costs, it is desirable to deploy systems that can operate fully independently. For a reconnaissance mission, this requires a planning of the complete mission before the drone leaves for enemy territory. This is a so-called Single-Hop Routing Protocol, as the the planning is updated in only one ‘hop’ [15]. The setting of such a mission can be stated as follows: starting from a secure base, multiple surveillance locations need to be safely reached and the acquired information has to be brought back to the base. There are three possible ways of bringing back information. At every surveillance location, there is the possibility of transmitting information back to the base camp. Hence the first option is that the drone travels to a surveillance location, gathers the info and immediately sends it back to the base. Secondly, the drone could also store the information and go to the next surveillance location. After having obtained the information there, the totality of info could then be transmitted together at that location. Thirdly, the drone could also store information and return to the base. In this case the information is physically obtained back from the drone. However, each action in the mission carries the risk of detection – the drone could be spotted during flight, or transmissions might be intercepted. Both ways of detection give away the position of the drone, ending the mission abruptly.
@@
This paper investigates how to find the optimal strategy of these reconnaissance missions. Such an optimal strategy consists of two elements: both the route and the send strategy have to be optimal to maximize the amount of retrieved information. Hence the specific questions for which an answer is proposed in this research are:
  • In which order should the different locations be travelled to?
@@
  • Where is it beneficial to make a transmission, and where is it better to hold on to the gathered information?
1.1 Related Work
Mission planning for UAV reconnaissance has been studied extensively in literature, as there are various factors to take into account. These factors can be subdivided into three categories: modelling of means, modelling of terrain and modelling of threat [22]. Every such category yields a different objective in the optimisation and a different way of approaching the problem.
Modelling of means encompasses all specific and intrinsic vehicle characteristics. A UAV has different features than a UGV, and hence needs to be modelled differently. Evidently, UAVs are able to move up and down, giving an additional dimension to the routing problem [22]. Furthermore, a small UAV could require a recharge during the mission, maybe even on a mobile charging station. This scenario was studied and modelled by Yu et al. [25]. But also other types of missions combine both types of unmanned vehicles; see Lu et al. [17] for a review. Modelling of means also comprises of the specific goals besides surveillance of the mission. Sajid et al. [20] propose a model for missions which also contain a delivery component.
Secondly, modelling of terrain comprises of all terrain elements the vehicle can encounter during the mission. Especially UGVs are constrained by the terrain. Roberge, Tarbouchi and Labonté [19] have divided the terrain into small square cells which represent areas of similar altitude. They then study the performance of genetic algorithms for routing across such a grid, taking into account fuel constraints coming from uphill and downhill movements. Fink et al. [8] have proposed an adaptation of Dijkstra’s algorithm for multi-objective routing in three-dimensional, mountainous terrain. Interestingly, they do so for routing a rover on Mars. But also when considering UAVs, the terrain plays a role. There might be altitude constraints, or constraints related to the angle of ascent and descent, especially when multiple UAVs are involved and collisions should be avoided; see Jia et al. [14] for an approach of this scenario. Furthermore, Albert, Leira and Imsland [1] considered the problem where multiple UAVs surveil moving targets. They propose an Mixed Integer Linear Program to model this situation and do a case study based on moving icebergs.
Lastly, modelling of threats includes all elements with an incentive to jeopardize the mission, e.g. radars and enemy lookout posts. Alotaibi et al. [2] model these threats as a cost function on the edges of a graph based on the distance to hostile waypoints. Dasdemir, Köksalan and Öztürk [5] consider a continuous terrain with circular radar zones. Another possible threat is jamming, making it impossible to send any information back to the basis. Han et al. [10] propose a reinforcement learning approach to deal with an uncertain jamming state when routing. There are multiple ways to deal with threats – proactive or reactive – which require different algorithms; see also Quadt et al. [18] for a recent and extensive survey.
1.2 Our Contribution
The main contribution of this paper is considering a novel characteristic in the model of means which implies a new threat to a reconnaissance mission. In a classical setting, it is assumed that the unmanned vehicle physically brings gathered information back to the basis, or the risk of detection when transmitting this info is not considered. In this work, we consider what happens when we do take into account the ability of transmission and the threat of interception. Hence, we assume that, based on the vehicle characteristics, modelling of the terrain, and similar techniques as in Alotaibi et al. [2], we have been given a graph with a cost function on the edges corresponding to the survival probability to cross that edge. But now we also add a cost function to the vertices, corresponding to surviving a potential transmission at that location. This added threat requires a new model and new solutions, which are presented below.
@@ 2.2 NP-completeness
The problem lies in NP. Let $G^{\prime}=(V^{\prime},E^{\prime})$ be an undirected graph such that $0,t\in V^{\prime}$. The problem of deciding whether there is a Hamiltonian path from 00 to $t$ in $G^{\prime}$ is NP-complete, see e.g. Schrijver [21, Corollary 8.11b]. Now construct the graph $G=(V,E)$ by adding a vertex $u$ to $G^{\prime}$, connecting it only to $t$. Formally, $V=V^{\prime}\cup\{u\}$ and $E=E^{\prime}\cup\{\{t,u\}\}$. Then $G$ contains a Hamiltonian path with starting point 00 if and only if $G^{\prime}$ contains a Hamiltonian path from 00 to $t$. ∎
@@ 3.1 Moving the drone
The route of the drone will be modelled by the decision variables
@@ 4.1.1 Initialization
To initialize the algorithm, a list of random routes is generated. These routes are constructed by making a random walk of random length through the graph, starting from the base. Just as in the Linear Program, where the value of $T_{\text{max}}$ had to be guessed, the upper bound $L_{\max}$ on the length of the walk also has to be chosen here. Similarly, we also choose a lower bound $L_{\min}$. Then we sample the actual length of the random walk uniformly at random from the set $\{L_{\min},L_{\min}+1,\dots,L_{\max}-1,L_{\max}\}$. Secondly, we repeatedly pick the next vertex in the walk uniformly at random from the vertices adjacent to the current position. If the last vertex of the walk is not the base vertex, the shortest path between the end of the walk and the base is added to the walk, using Dijkstra’s algorithm [6]. As a rule of thumb, when considering a graph on $n$ vertices, we picked $L_{\min}$ to be approximately equal to $n-1$ and $L_{\max}$ to $n+100$. This might seem like a pretty high upper bound, especially because realistically, reconnaissance mission do not include more than 10 surveillance locations. But even though these longer routes are not necessarily better than the shorter ones – at some point all information has been transmitted and making an additional loop doesn’t change the expected value – the final part of a route with useless and unnecessary walks can become useful after a crossover.
@@ 5.1 Adaptation of the Genetic Algorithm
Also the mutations have the same flavour as in the single drone case. First of all, there is the added random walk. This is the same mutation as described for the single drone: one of the routes is picked and a random walk is added somewhere in the walk. This continues to be a useful mutation to increase the variation in the genetic program. But it slows down the algorithm, because it requires a local search for the optimal send strategy in terms of the expected value for multiple drones. Therefore this mutation is only actualized with a low probability. If a generation consists of $\kappa$ strategies, the mutation probability is set at 2/$\kappa$.